OpenStudy (anonymous):

Multiply and simplify by factoring. ^3sqrty^10 3sqrt16y^11=

OpenStudy (anonymous):

\[\sqrt[3]{y^{10}} \sqrt[3]{16y^{11}}\]

OpenStudy (anonymous):

Is that it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. Easiest way is to just write everything out so you can see. y^10 = (y*y*y) * (y*y*y) * (y*y*y) * y

OpenStudy (anonymous):

Any idea why I grouped those 10 that way?

OpenStudy (anonymous):

because they cancel out?

OpenStudy (anonymous):

Well, since you are taking the cube root, I wanted to group them in groups of 3. For each group of 3, you can pull that outside the radical.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So that means I can pull out y^3

OpenStudy (anonymous):

\[y^3\sqrt[3]{y}\]

OpenStudy (anonymous):

16 = (2*2*2) * 2 * 2

OpenStudy (anonymous):

and y^11 = (y*y*y) * (y*y*y) * (y*y*y) * y * y

OpenStudy (anonymous):

So we can pull out 2y^3

OpenStudy (anonymous):

But now we have (y^3)(2y^3) ^3sqrt(y) ^sqrt(y^2) Combine those last two and we get ^3sqrt(y^3) and that means another y comes out so we are left with (y^3)(2y^3)(y)

OpenStudy (anonymous):

Which is 2y^7

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

2y^7 ^3sqrt(4)

OpenStudy (anonymous):

here is another way for the cognoscenti. 3 (the index) goes in to 10 (the exponent) 3 times with a remainder of 1. so \[y^3\] comes out and one y stays in

OpenStudy (anonymous):

Yeah that is a lot easier. :) Just want to make sure she understands why that is the case.

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

The other thing too is that at the beginning we should have combined everything to get 16y^21 underneath the radical.

OpenStudy (anonymous):

It is then easy to see there are 7 groups of (y*y*y)